# Universal Deformations and Inhomogeneities in Isotropic Cauchy Elasticity

### v1

@@ 5 Conclusions
The existing studies of universal deformations have been restricted to hyperelasticity, or more generally, to materials that have an underlying energy function. In this paper, we extended the analysis of universal deformations to Cauchy elastic solids, which in general, do not have an energy function. We considered both compressible and incompressible inhomogeneous isotropic Cauchy elasticity. As hyperelasticity is a proper subset of Cauchy elasticity, one would expect the set of universal deformations of Cauchy elasticity to be a proper subset of that of hyperelasticity. We proved the unexpected result that the sets of universal deformations of isotropic Cauchy elasticity and isotropic hyperelasticity are identical, in both the compressible and incompressible cases. We also proved that their corresponding universal inhomogeneities are identical as well.

### v2

@@ 5 Conclusions
The existing studies of universal deformations have been restricted to hyperelasticity, or more generally, to materials that have an underlying energy function. In this paper, we extended the analysis of universal deformations to Cauchy elastic solids, which in general, do not have an energy function. We considered both compressible and incompressible inhomogeneous isotropic Cauchy elasticity. As hyperelasticity is a proper subset of Cauchy elasticity, one would expect the set of universal deformations of Cauchy elasticity to be a proper subset of that of hyperelasticity. We proved the somewhat unexpected result that the sets of universal deformations of isotropic Cauchy elasticity and isotropic hyperelasticity are identical, in both the compressible and incompressible cases. We also proved that their corresponding universal inhomogeneities are identical as well.